# Electrocardiogram Instruction Tuning for Report Generation

### v2

@@ Appendix A Hyper-parameters of ECG Instruction Tuning
In this study, we concentrate on the implementation of the Low-Rank Adaptation (LoRA) (Hu et al., 2021) technique for efficient fine-tuning, specifically applied to ECG instruction tuning. As detailed in Table A provided, we utilize mixed precision at bf16 for enhanced computational efficiency. Our models undergo instruction tuning over 5 epochs, with LoRA parameters set at an alpha of 64 and a rank of 128, accompanied by a dropout rate of 0.1. The total batch size is 64 with a gradient accumulation factor of 2. The maximum sequence length is constrained to 256 tokens. Additionally, we adopt a learning rate with 2e-5 for GPT-NeoX, and 1e-4 for the other models, optimized using the AdamW algorithm. The learning rate follows a linear schedule with a warm-up ratio of 0.03. We set the weight decay to 0.0.
@@
Moreover, as shown in Table A, we detail the ECG embedding dimensions for various language models, highlighting their approach to ECG data encoding. GPT2-Medium and GPT2-Large feature ECG dimensions of 64, while GPT-Neo, BLOOM, OPT, LLaMA-1, Mistral, LLaMA-2, and Mistral-Instruct use a dimension of 128. GPT-NeoX employs a dimension of 96, and GPT-J notably uses the largest dimension of 256. These dimensions, reflecting each model’s head dimension design, illustrate diverse strategies in ECG data processing across different models.
Appendix B Visualization of Generated ECG Report Samples
@@
As illustrated in Figures 9,  10, and  11 we have visualized the report samples generated by LLaMA-1, LLaMA-2, and Mistral-Instruct. The samples are presented in blue font to highlight the key information that aligns with the ground truth. The visualization demonstrates that all three selected models are capable of capturing the essential patterns of ECG signals and generating accurate reports. This underscores the efficacy of our proposed MEIT framework, which proves to be adaptable to the majority of open-source LLMs. It effectively learns the correct clinical semantics of ECG signals, thereby enabling the generation of corresponding reports.

### v3

@@ A.2 Hyper-parameters of ECG Instruction Tuning
In this study, we implement the Low-Rank Adaptation (LoRA) [16] technique for efficient fine-tuning, specifically applied to ECG instruction tuning. As detailed in Table A.2 provided, we utilize mixed precision at bf16 for enhanced computational efficiency. Our models undergo instruction tuning over 5 epochs, with LoRA parameters set at an alpha of 64 and a rank of 128, accompanied by a dropout rate of 0.1. The total batch size is 64, with a gradient accumulation factor of 2. The maximum sequence length is constrained to 256 tokens. Additionally, we adopt a learning rate with 2e-5 for GPT-NeoX and 1e-4 for the other models, optimized using the AdamW algorithm. The learning rate follows a linear schedule with a warm-up ratio of 0.03. We set the weight decay to 0.0.
@@
Moreover, as shown in Table A.2, we detail the ECG embedding dimensions for various language models, highlighting their approach to ECG data encoding. GPT2-Medium and GPT2-Large feature ECG dimensions 64, while GPT-Neo, BLOOM, OPT, LLaMA-1, Mistral, LLaMA-2, and Mistral-Instruct use a dimension of 128. GPT-NeoX employs a dimension of 96, and GPT-J notably uses the largest dimension of 256. These dimensions, reflecting each model’s head dimension design, illustrate diverse strategies in ECG data processing across different models.
A.3 More Details of ECG Encoder
Projection Layer For the design of the projection layer within the ECG encoder, we adopt a non-linear approach similar to CLIP [56] and Med-UniC [57]. Specifically, in our experiments, we employ two consecutive linear layers, each followed by BatchNorm1d [^4] https://pytorch.org/docs/stable/generated/torch.nn.BatchNorm1d.html. Besides, ReLU serves as the activation function between the two linear layers. The default settings for input and hidden layers dimensions are set to 2048 in our experiment.
Parameter Size Analysis To demonstrate the ECG encoder’s lightweight design, we analyzed its trainable parameters during instruction tuning and total parameters during inference, using the LLaMA-1 7B model for illustration (Table 7). The analysis reveals the ECG encoder’s trainable parameters are substantially fewer than those of the LoRA adapter in the LLM backbone during instruction tuning, and its parameter share of the overall framework is minimal for inference, underscoring its efficiency.
Table 7: Parameter Comparison of ECG encoder and LLM backbone. We use LLaMA-1 7B as an example.
Module Trainable Params Inference Params
LLM backbone 159M 6.90B
ECG encoder 20.4M 20.4M
A.4 Further Analysis of Generated Prompts
In the ECG instruction data curation, we manually created 32 prompt examples, as illustrated in Section 3.1. To increase the diversity of our samples, we employed GPT-4 to rephrase these manually designed prompts, generating a larger pool of prompt examples. These generated examples were randomly sampled and paired with ECG-text pairs to compile the ECG instruction dataset. In this section, We compare the experiment’s effects using 128, 256, and 512-generated samples, respectively. Table 8 shows the corresponding results with different dimensions. When the number is 256, it can achieve better results in most experimental settings. Hence, we take 256 generated samples as our default setting during the instruction tuning and inference.
Table 8: Performance comparison of different numbers of generated prompt samples. We evaluate them on the MIMIC-IV-ECG dataset, using BLEU-4, METEOR, ROUGE-L, and CIDEr-D metrics. We take LLaMA-1 7B as the LLM backbone here. heavy teal color indicates the highest results.
Prompt Nums BLEU-4 METEOR ROUGE-L CIDEr-D
128 0.541 0.756 0.718 5.15
256 0.543 0.761 0.724 5.26
512 0.538 0.754 0.732 5.03
A.5 Visualization of Generated ECG Report Samples
@@
As illustrated in Figures 9,  10, and  11 we have visualized the report samples generated by LLaMA-1, LLaMA-2, and Mistral-Instruct. The samples are presented in blue font to highlight the key information that aligns with the ground truth. The visualization demonstrates that all three selected models can capture the essential patterns of ECG signals and generate accurate reports. This underscores the efficacy of our proposed MEIT framework, which is adaptable to most open-source LLMs. It effectively learns the correct clinical semantics of ECG signals, thereby enabling the generation of corresponding reports.